OpenStudy (anonymous):

f(x)=x^2-x^12+1 odd even neither

OpenStudy (anonymous):

It's shifted to the right of the y-axis, so neither.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Do you know both the function-evaluation definition and the graphical definition of odd vs. even?

OpenStudy (unklerhaukus):

it is shifted up the y-axis, not to the right , along the x-axis @CliffSedge

OpenStudy (anonymous):

Oh, woops, must have had my eyes crossed, I read the middle term as 12x instead of x^12! X-D

OpenStudy (anonymous):

Yes, Unkle is correct, all three terms are of even degree.

OpenStudy (unklerhaukus):

\[f(x)=x^2-x^{12}+1\]\[\downarrow\]\[f(x)=x^2-x^{12}+x^0 \]

OpenStudy (anonymous):

\[f(-x)=f(x) \rightarrow even.\]

OpenStudy (unklerhaukus):

\[f(-x)=(-x)^2-(-x)^{12}+(-x)^0=x^2-x^{12}+x^0=f(x)\] \[\therefore \text{even}\]

OpenStudy (anonymous):

Got all that, @avatar2012 ?